An air table has two pucks; one with a mass of 0.50kg and the other with a mass of 1.50kg.  Initially the smaller puck is moving at 0.40m/s, 0° while the larger puck is moving at 0.40m/s, 180°.  After the collision, the smaller puck is moving at 0.79m/s, 196°.  What is the velocity (magnitude and direction) of the larger puck after the collision?



1
Expert's answer
2021-12-06T16:41:49-0500

"m_1= 0.5kg"

"v_1= 0.4\frac{m}{s}"

"\alpha_1= 0\degree"

"\vec v_1(v_1\cos \alpha_1,v_1\sin \alpha_1)"

"\vec v_1(0.4;0)"

"m_2= 1.5kg"

"v_2= 0.4\frac{m}{s}"

"\alpha_2= 180\degree"

"\vec v_2(v_2\cos \alpha_2,v_1\sin \alpha_2)"

"\vec v_2(-0.4;0)"

"u_1 = 0.79\frac{m}{s}"

"\alpha'_1= 196\degree"

"\vec u_1(u_1\cos \alpha'_1,u_1\sin \alpha'_1)"

"\vec u_1(-0,76;-0.22)"

"\text{According to the law of momentum:}"

"m_1\vec v_1+ m_2\vec v_2=m_1\vec u_1+m_2\vec u_2"

"0.5*(0.4;0)+1.5*(-0.4;0)= 0.5*(-0.76;-0.22)+1.5*(u_{2x};u_{2y})"

"\vec u_2(-0.013;0.073)"

"u_2=\sqrt{0.013^2+0.073^2}=0.074\frac{m}{s}"

"\cos \alpha'_2=\frac{-0.013}{0.074}=-0.17568"

"\alpha'_2= 100.11\degree"


"\text{Answer: }u_2=0.074\frac{m}{s};\alpha'_2= 100.11\degree"